OpenStudy (leozap1):

Is anyone here good at multiplying binomials that will explain to me how to do it I know for the most part I just have a few questions on it

OpenStudy (anonymous):

well what exactly is your question?

OpenStudy (leozap1):

Well I don't have a question in general I am working on them in school and I don't fully understand them I know how to get the first & last term but not the second/middle term

OpenStudy (leozap1):

I use FOIL to get the answer

OpenStudy (anonymous):

yeh FOIL is the easiest way BUT honestly it would be better to ask your teacher to explain it to you because i dont want to give false information xD

OpenStudy (leozap1):

I go to a online school and it would take a good bit for my teacher to reply to me and most likely she will ask me to go a live lesson that would be like a week away from now

OpenStudy (anonymous):

well ummmmmm idrk xD I'm just waking up and at school right now. maybe later i can help? again sorry

OpenStudy (leozap1):

Ok thanks anyway

OpenStudy (atlas):

(2+3)(4+5) = 2(4+5) +3(4+5) do u get this?

OpenStudy (anonymous):

oh gawd, numbers O.o

OpenStudy (atlas):

this is simple distributive law of multiplication over addition

OpenStudy (atlas):

I am going to come to binomials too :P

OpenStudy (leozap1):

@atlas I already know how to do that the kind of questions I am having trouble with is like (2x-3)^2 = (2x-3)(2x-3) I know how to get the first and last terms but not the second one it would be like 4z^2 - (something) + 9

OpenStudy (atlas):

oh it is very similar to what i wrote (2x-3)(2x-3) = 2x(2x-3) - 3(2x-3)

OpenStudy (atlas):

See if you get it now :P

OpenStudy (atlas):

There is another formulae for doing it but I hope you will understand it better if you do it this way

OpenStudy (leozap1):

No I don't sorry and to answer the first question you ask would be 8+10+12+15=45

OpenStudy (atlas):

oh.......i was not asking you the answer of the expression do you understand why i wrote (2+3)(4+5) as 2(4+5) +3(4+5)

OpenStudy (leozap1):

I was taught to use FOIL but I know how you got that

OpenStudy (atlas):

ok same way (2x-3)(2x-3) can be written as 2x(2x-3) -3(2x-3) Do u get this?

OpenStudy (leozap1):

Yes I do you are multiplying the second binomial by each term in the first binomial

OpenStudy (atlas):

Exactly! You can also understand it this way. I have to multiply (2x-3) with (2x-3). Now that is lot of work!! SO I multiply first 2x with (2x-3) and then -3 with (2x-3)

OpenStudy (atlas):

(2x-3)(2x-3) = 2x(2x-3) -3(2x-3) = (2x*2x) - (2x*3) - (3*2x) - (3*-3)

OpenStudy (atlas):

Again I have to multilpy 2x with (2x-3) and that is lot of work again........>So I multiply 2x with 2x first and then 2x with -3

OpenStudy (atlas):

Add them up this way now........Take care of the signs and you will get your answer

OpenStudy (leozap1):

Ok so for the first term in the solution I take the first term in the problem and multiply it by its self for the second term in the solution I take the second term in the problem and multiply it (How you got that) I don't know how you got that...

OpenStudy (atlas):

Ah......it is called distributive law of multiplication over addition It is the same thing I did earlier and you said you got it (2+3)(4+5) = 2(4+5) +3(4+5)

OpenStudy (leozap1):

Ok but how did you get these binomials? (2x*3) - (3*2x)